OpenStudy (anonymous):

Express the following function, F(x) as a composition of two functions f and g .

OpenStudy (anonymous):

\[f(x)=x^2/(x^2+4)\]

OpenStudy (anonymous):

https://www.surveymonkey.com/s/YRBHK56

OpenStudy (anonymous):

f(x)=x^2-4 a=1,b=0,c=-4 .. x1= (0+sqrt(0+16))/2 x1=2 x2= (0-sqrt(0+16))/2 x2= -2

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

wc